vectors are represented along a hexagon OPQRST in order. if op=A AND PQ=B.FIND THE VECTORS ALONG OTHER SIDES OF THE HEXAGON IN ORDER

Grade:12

1 Answers

bramashini
5 years ago
let’s assume in the hexagon extended OP and RQ will meet at U. Hence, PU=OP=A 
PQU is equlateral triangle.
QR=OT= UQ = UP+PQ = -A+B = B-A (COMMUTATIVE LAW)
RS = OP =A SO, SR = -A 
ST = QP = -B
TO = QU = A-B
These are the vectors acting along the sides of the hexagon (i assumed hexagon is regular one therefore  PQU IS EQULATERAL TRIANGLE  the one of the angles of hexagon = 120 degrees )